OpenStudy (anonymous):

Which of the following is a solution to the equation log 6 x + log 6 (x + 1) = 1 ?

OpenStudy (anonymous):

log6 x - log6 (x-1) =1 log6 (x/(x-1) = 1 6 to the power of one = x/(x-1) 6 = x/(x-1) 6x - 6 = x 5x = 6 x = 6/5 = 1.2

OpenStudy (anonymous):

thats not what I put in I said log 6 x **+ log 6 (x **+ 1) = 1

OpenStudy (anonymous):

unless that is the same @Jenna_Da_Nerd ..

OpenStudy (anonymous):

oops looked at it wrong

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

i saw this on another open study question already with the (-)'s @Jenna_Da_Nerd thanks for your help so far though

OpenStudy (anonymous):

this isnt the answer though

OpenStudy (anonymous):

I know.. If you could help with this one I would appreciate it(:

OpenStudy (anonymous):

Subtracting logs means divide so we can combine on the right, for example: multiply both sides by x x(x+1) = 6 arrange as a quadratic equation x^2 + x - 6 = 0 Factors to (x+3)(x-2) Two solutions x = -3 x = +2, this is the only solution, (we can't find the log of a neg number) Check this on a calc: enter: log(2+1) results: .477

OpenStudy (anonymous):

just adding that the 1 is log6(6)

OpenStudy (anonymous):

thats just an example to help u

OpenStudy (anonymous):